OpenStudy (anonymous):

If x,y, and z are positive numbers and 2x=3y=4z, then the value of x+y+z is how many times the value of x?

OpenStudy (anonymous):

2x=3y, so y=2x/3 2x=4z, so z=x/2 x+y+z=(1+2/3+1/2)*x=13x/6 13/6

OpenStudy (anonymous):

ok thank you